OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 5+3x=sin(xy)^2

OpenStudy (anonymous):

d/dx(5+3x) = d/dx(sin(xy)^2 so the left side is pretty obvious...I'm guessing its the RHS that's giving you trouble.

OpenStudy (anonymous):

\[d/dx \sin (xy)^2 = d/dx sinu *d/dx (u) \] if we use the chain rule and (yx)^2=u

OpenStudy (anonymous):

i got cos(xy)^2*2xy* 1*y*y' for that part but idk if its right

OpenStudy (anonymous):

so d/dx (u) is where implicit differentiation comes in. You'll have to use two rules: the product rule of differentiation (d/dt (xy) = (x)'*y + x*(y)' ) and the implicit differentiation rule.

OpenStudy (anonymous):

i no how to do the product rule, its just the implicit differientiation rule that confuses me

OpenStudy (amistre64):

rules for implicit are exactly the same for explicit; only thing is you want to keep your x' and y' til the end then factor out your y'. your x'[dx/dx] will of course equal 1

OpenStudy (anonymous):

so when u differenciate the y wat does that become? y*y'?

OpenStudy (amistre64):

yes, y becomes y' 6y^2 becomes 12y y'

OpenStudy (anonymous):

oooooooooooooooooooo!!!!

OpenStudy (amistre64):

xy becomes: x'y + xy' not that it helps here....maybe

OpenStudy (amistre64):

...but it might lol

OpenStudy (anonymous):

hmmm...I was thinking it would, since d(xy)/dx) = y+x*dy/dx?

OpenStudy (amistre64):

2(sin(xy)) cos(xy) (x'y + xy') ??

OpenStudy (anonymous):

so when i differenciate the whole thing it becomes 0+3=cos(xy)^2 * 2xy* (1)(y)(y')?

OpenStudy (anonymous):

or not...

OpenStudy (amistre64):

u^2 Du ; u = sin(t); t = xy

OpenStudy (amistre64):

2sin(xy) cos(xy) (x'y + xy') does that help out?

OpenStudy (anonymous):

yaa so thats for the right side only right

OpenStudy (anonymous):

so its a chain rule within a chain rule?

OpenStudy (amistre64):

yep; chain a chain lol

OpenStudy (amistre64):

D(u) D(sin(t)) D(t) is what I get from it

OpenStudy (anonymous):

ooooooo ok i get it!

OpenStudy (anonymous):

so now i solve for y'

OpenStudy (amistre64):

yep, x' = 1 so we can ignore those; and solve for y'

OpenStudy (anonymous):

do i divide 2sinxy*cos(xy) on both sides?

OpenStudy (amistre64):

its a good start :)

OpenStudy (amistre64):

then subtract a "y" and divide it all again by "x"

OpenStudy (anonymous):

so i got 3-y/2sin(xy)*cos (xy)*x

OpenStudy (anonymous):

is that correct?

OpenStudy (amistre64):

thats a little rough, not quite it....

OpenStudy (anonymous):

awwww!!! lol

OpenStudy (anonymous):

wat did i do wrong?

OpenStudy (amistre64):

3 y ---------------- - -- x (2sin(xy)cos(xy)) x

OpenStudy (anonymous):

where did the x under the 3 come from?

OpenStudy (amistre64):

lets say B = (2sin(xy)cos(xy)) to clean this up... 3/B = y + xy' (3/B) -y = xy' (3/B)/x - y/x = y' 3/Bx - y/x = y'

OpenStudy (anonymous):

oo i turn the x into a 1

OpenStudy (amistre64):

x' = dx/dx = 1 like any good fracation does :)

OpenStudy (anonymous):

ooooooooo ok i see wat i did wrong

OpenStudy (amistre64):

you got 3-y/Bx

OpenStudy (anonymous):

omg im gonna do exactly wat u did for the test.. change that whole part to a B so i wont make a mistake

OpenStudy (amistre64):

lol ..... it is easier on the eyes fer sure :)

OpenStudy (anonymous):

THANK U SOO MUCH!!!

OpenStudy (anonymous):

UR A LIFE SAVER!!